A particle of mass m moves with momentum of magnitude p. (a) Show that the kinetic energy of the particle is K = p^2 / 2m. (b) An object has a kinetic energy of 275 J and a momentum of magnitude 25.0 kgm/s. Find the speed and mass of the object.
1
Expert's answer
2020-11-25T07:14:51-0500

Since kinetic energy is given by "K. E. =\frac{1}{2}mv^2"


Momentum is given by, "P=mv"

"\implies v=P/m"


Then, I can write it as,

"K. E. =\frac{1}{2}mv^2=\frac{1}{2}m(p/m)^2= \frac{p^2}{2m}"



Given, kinetic energy, "K. E. =\frac{1}{2}mv^2 =275J"


Momentum, "P =mv=25.0 kgm\/s"


Dividing both, "\frac{mv^2}{2mv}=\frac{275}{25}"

"v=2\times 11=22m/s"


Using the relationship which we have derived,

"K. E. = \frac{p^2}{2m}"


"275=\frac{25^2}{2m}"


"m =1.14 Kg"